OpenStudy (anonymous):

How would you work out this equation? could give a detailed explaination? these problems are driving me nuts. 2.1x + 45.2 = 3.2 - 8.4x

OpenStudy (saifoo.khan):

ADD LIKE TERMS ALWAYS!

OpenStudy (vijay):

2.1x +8.4x = 3.2 -45.2 10.5x = -42 x= -4

OpenStudy (anonymous):

vijay can you show me the steps you used to get -4?